Chiara's note:
BSRTMB Beam 1 - operation feedbacks:
- ask availability of the alignment team for a measurement of the tank during TS1 --> Daniele/William
BSRTMBs Beam 1 & 2, HL integration:
- check of the beam aperture from D4 to the BSRTMB tank --> any aperture smaller than our tank is a potential restriction to light --> William
- Centre the tanks longitudinally wrt to the existing carrotage, optimising its position wrt to space availability for the optical line --> William + Chiara + HL Integration
- keep the enlarged aperture of the vacuum chambers to a DN100
- the vacuum chambers will be NEG coated, any need of aC coating to reduce the signal from reflected photons? --> Chiara to check with Marton the possibility to use SynRad. Note: I've contacted him, they schedule a training in september. For a 1 to 1 training he wanted to check with his boss.. I doubt he is available for helping with the study but I'll keep asking. In the meantime I said I'll go to the training.
- Integrate a second vacuum chamber with a port, same as the BTVSI, as close as possible to the D4, this is used to mimick the light during installation and align the BSRTMBs. Integration to be optimised wrt to the shottky on B2. The swap is not absolutely crucial for the light, but the integration of a new tank is. --> William + Chiara + HL Integration
- Study of the tunnel vibrations is crucial to understand the stability of the optical line. Contact M. Guinchard for equipping a dummy line to be installed in YETS in the carrotage. --> Chiara, William
- Check with transport the envelop available considering option 3 and 4 of the optical line (so below and above the TIM) --> Chiara/William
- Check James is still happy with rotating the racks 90deg ---> Chiara
- optical line specs: no need of vacuum, but light and air tight with two viewports. As much as possible vibration free. Space reservation volume 30cm, depending on the cable occupancy of the existing carrotage.
Stephane's note:
STATUS OF COMMISSIONING
BSRT:
- As probe were found disconnected, Daniele added a probe status in the watchdog BSRT.
- Beam 1 OK
- Beam 2 signal issue à beam tail + large (not on focal point)
- BSRT/H splitting checked OK (give the same!)
- Access Tuesday 15/04 à check any potential reflection, put some margin on L2 stepper
BSRL:
- B2 OK
- B1: signal very low (~100x less than expected/last year)
BSRA: No change
BSRAM: not before Clara comes back (2 weeks)
BSRH: Functionality OK
KICK-OFF SECONDARY EXTRACTION LINES
In addition of Chiara’s clear minutes and actions.
Quantifying jitter effect: simulation tool by Daniele
- 50m optical line, vibrating 1 (out of 5? or all 5?) mirror by 1urad ends up to 150um PSF
- Would be nice information to check in lab jitter of image on a 30(?)m long optical line in air (with/without covers ?) Stephane/Jan/Daniele
BI STEPPER SITUATION
- Need stepper from CEM for BSRTMB mirrors (interlock + readback)
- Need ‘standard’ stepper controller (without feedback) à 2x30 = 60 for the BSRTMBs
- Options:
- CEM (Sambuca). INFO: Sambuca version 2 is more compact (from racks to a DIOT crate format for 6 channels): Cost ? OK for LS3?
- BI/PM (VFC/GEFE): no spares.
- BI/PM (Peter & David’s version based on SoC & Ethernet): needs to be finalized and deployed. Cost? OK for LS3 ?
- Need a clear statement from BI hierarchy. Same question on the table for the >40 or 50 new stepper need for AWAKE à Federico to discuss with Tibo ??
There are minutes attached to this event.
Show them.